solve the equation.
/2s-3/=/s+7/
Also could you please explain steps. I would greatly appreciate it.

Answer by josgarithmetic(39617)

Without a detailed explanation, either 2s-3=s+7 or 2s-3=-s-7.

The first condition gives
2s-s-3=7
s-3=7
s=10.

The second condition gives
2s+s-3=-7
3s-3=-7
3s=-7+3
3s=-1
s=-1/3